# Vision Transformer with Sparse Scan Prior

### v1

@@ Abstract
In recent years, Transformers have achieved remarkable progress in computer vision tasks. However, their global modeling often comes with substantial computational overhead, in stark contrast to the human eye’s efficient information processing. Inspired by the human eye’s sparse scanning mechanism, we propose a Sparse Scan Self-Attention mechanism ($\rm{S}^{3}\rm{A}$). This mechanism predefines a series of Anchors of Interest for each token and employs local attention to efficiently model the spatial information around these anchors, avoiding redundant global modeling and excessive focus on local information. This approach mirrors the human eye’s functionality and significantly reduces the computational load of vision models. Building on $\rm{S}^{3}\rm{A}$, we introduce the Sparse Scan Vision Transformer (SSViT). Extensive experiments demonstrate the outstanding performance of SSViT across a variety of tasks. Specifically, on ImageNet classification, without additional supervision or training data, SSViT achieves top-1 accuracies of 84.4%/85.7% with 4.4G/18.2G FLOPs. SSViT also excels in downstream tasks such as object detection, instance segmentation, and semantic segmentation. Its robustness is further validated across diverse datasets. Code will be available at https://github.com/qhfan/SSViT.
@@ 3.2 Sparse Scan Self-Attention
$\centering\begin{split}\mathbb{A}_{i,j}=&\{X_{m,n}|m=i+k_{h}\times S_{h},n=j+k_{w}\times S_{w}\},\\ &k_{h}\in(-\lfloor N_{h}/2\rfloor,\lfloor N_{h}/2\rfloor),k_{h}\in\mathbb{Z},\\ &k_{w}\in(-\lfloor N_{w}/2\rfloor,\lfloor N_{w}/2\rfloor),k_{w}\in\mathbb{Z}.\\ \end{split}\@add@centering$ (2)
@@ 4.1 ImageNet Classification
Cost Model Parmas (M) FLOPs (G) Top1-acc (%)
tiny model $\sim 2.5$G QuadTree-B-b1 [57] 14 2.3 80.0
@@
small model $\sim 4.5$G Swin-T [40] 29 4.5 81.3
@@
Cost Model Parmas (M) FLOPs (G) Top1-acc (%)
base model $\sim 9.0$G ConvNeXt-S [41] 50 8.7 83.1
@@
large model $\sim 18.0$G DeiT-B [59] 86 17.5 81.8
@@
Model Parmas (M) FLOPs (G) Throughput (imgs/s) Top1-acc (%)
@@
SS-Swin-B 88 15.7 538 85.1(+1.8)
Model Parmas (M) FLOPs (G) Throughput (imgs/s) Top1-acc (%)
@@ 4.2 Object Detection and Instance Segmentation
We utilize MMDetection [5] to implement Mask-RCNN [24], Cascade Mask R-CNN [2], and RetinaNet [37] for evaluating our SSViT. For Mask R-CNN and Cascade Mask R-CNN, we adhere to the commonly used “3×+3$\times$+3 × + MS” setting, and for Mask R-CNN and RetinaNet, we apply the “1×1$\times$1 ×” setting. Following [40], during training, we resize the images such that the shorter side is 800 pixels while keeping the longer side within 1333 pixels. We employ the AdamW optimizer for model optimization.
Backbone Params (M) FLOPs (G) Mask R-CNN 3×3$\times$3 ×+MS
@@
Backbone Params (M) FLOPs (G) Cascade Mask R-CNN 3×3$\times$3 ×+MS
@@
Table 3: Comparison with other backbones using "$3\times+\mathrm{MS}$“ schedule on COCO.
@@ Results.
Tab. 3 and Tab. 4 present the performance of SSViT across different detection frameworks. The results highlight that SSViT consistently outperforms its counterparts in all comparisons. Under the “3×+3$\times$+3 × +MS” schedule, SSViT surpasses the recent SMT, achieving a +2.2 box AP and +2.0 mask AP improvement with the Mask R-CNN framework. For Cascade Mask R-CNN, SSViT still maintains a significant performance edge over SMT. Regarding the “1×1$\times$1 ×” schedule, SSViT exhibits remarkable performance. Specifically, SSViT-S attains an improvement of +2.2 box AP and +1.5 mask AP over InternImage-T within the Mask-RCNN framework.
Backbone Params (M) FLOPs (G) Mask R-CNN 1×1$\times$1 × Params (M) FLOPs (G) RetinaNet 1×1$\times$1 ×
@@
Table 4: Comparison to other backbones using "1×1$\times$1 ד schedule on COCO.
@@ 5 Conclusion
SS-Swin-T [2, 2, 6, 2] [96, 192, 384, 768] [3, 6, 12, 24] 4 29 4.8
SS-Swin-S [2, 2, 18, 2] [96, 192, 384, 768] [3, 6, 12, 24] 4 50 8.9
SS-Swin-B [2, 2, 18, 2] [128, 256, 512, 1024] [4, 8, 16, 32] 4 88 15.7
SS-FasterViT-0 [2, 3, 6, 5] [64, 128, 256, 512] [–, –, 8, 16] 4 25 3.4
SS-FasterViT-1 [1, 3, 8, 5] [80, 160, 320, 640] [–, –, 8, 16] 4 41 5.4
SS-FasterViT-2 [3, 3, 8, 5] [96, 192, 384, 768] [–, –, 12, 24] 4 58 8.8
@@ Appendix E Code
18 self.qkv = nn.Conv2d(embed_dim, embed_dim*3, 1, bias=True)
19 self.out_proj = nn.Conv2d(embed_dim, embed_dim, 1, bias=True)

### v2

@@ Abstract
In recent years, Transformers have achieved remarkable progress in computer vision tasks. However, their global modeling often comes with substantial computational overhead, in stark contrast to the human eye’s efficient information processing. Inspired by the human eye’s sparse scanning mechanism, we propose a Sparse Scan Self-Attention mechanism ($\rm{S}^{3}\rm{A}$). This mechanism predefines a series of Anchors of Interest for each token and employs local attention to efficiently model the spatial information around these anchors, avoiding redundant global modeling and excessive focus on local information. This approach mirrors the human eye’s functionality and significantly reduces the computational load of vision models. Building on $\rm{S}^{3}\rm{A}$, we introduce the Sparse Scan Vision Transformer (SSViT). Extensive experiments demonstrate the outstanding performance of SSViT across a variety of tasks. Specifically, on ImageNet classification, without additional supervision or training data, SSViT achieves top-1 accuracies of 84.4%/85.7% with 4.4G/18.2G FLOPs. SSViT also excels in downstream tasks such as object detection, instance segmentation, and semantic segmentation. Its robustness is further validated across diverse datasets.
@@ 3.2 Sparse Scan Self-Attention
[…]
@@ 4.1 ImageNet Classification
Cost
Model
Parmas
FLOPs
Top1-acc
tiny model $\sim 2.5$G
QuadTree-B-b1 [57] 14 2.3 80.0
@@
small model $\sim 4.5$G
Swin-T [40] 29 4.5 81.3
@@
Cost
Model
Parmas
FLOPs
Top1-acc
base model $\sim 9.0$G
ConvNeXt-S [41] 50 8.7 83.1
@@
large model $\sim 18.0$G
DeiT-B [59] 86 17.5 81.8
@@
Model
Parmas
FLOPs
Throughput
(imgs/s)
Top1-acc
@@
$S$-Swin-B 88 15.7 538 85.1(+1.8)
Model
Parmas
FLOPs
Throughput
(imgs/s)
Top1-acc
@@ 4.2 Object Detection and Instance Segmentation
We utilize MMDetection [5] to implement Mask-RCNN [24], Cascade Mask R-CNN [2], and RetinaNet [37] for evaluating our SSViT. For Mask R-CNN and Cascade Mask R-CNN, we adhere to the commonly used $3\times+$ MS” setting, and for Mask R-CNN and RetinaNet, we apply the $1\times$” setting. Following [40], during training, we resize the images such that the shorter side is 800 pixels while keeping the longer side within 1333 pixels. We employ the AdamW optimizer for model optimization.
Backbone
Params
FLOPs
Mask R-CNN $3\times$+MS
@@
Backbone
Params
FLOPs
Cascade Mask R-CNN $3\times$+MS
@@
Table 3: Comparison with other backbones using $3\times+\mathrm{MS}$“ schedule on COCO.
@@ Results.
Tab. 3 and Tab. 4 present the performance of SSViT across different detection frameworks. The results highlight that SSViT consistently outperforms its counterparts in all comparisons. Under the $3\times+$MS” schedule, SSViT surpasses the recent SMT, achieving a +2.2 box AP and +2.0 mask AP improvement with the Mask R-CNN framework. For Cascade Mask R-CNN, SSViT still maintains a significant performance edge over SMT. Regarding the $1\times$” schedule, SSViT exhibits remarkable performance. Specifically, SSViT-S attains an improvement of +2.2 box AP and +1.5 mask AP over InternImage-T within the Mask-RCNN framework.
Backbone
Params
FLOPs
Mask R-CNN $1\times$
Params
FLOPs
RetinaNet $1\times$
@@
Table 4: Comparison to other backbones using $1\times$“ schedule on COCO.
@@ 5 Conclusion
$S$-Swin-T [2, 2, 6, 2] [96, 192, 384, 768] [3, 6, 12, 24] 4 29 4.8
$S$-Swin-S [2, 2, 18, 2] [96, 192, 384, 768] [3, 6, 12, 24] 4 50 8.9
$S$-Swin-B [2, 2, 18, 2] [128, 256, 512, 1024] [4, 8, 16, 32] 4 88 15.7
$S$-FasterViT-0 [2, 3, 6, 5] [64, 128, 256, 512] [–, –, 8, 16] 4 25 3.4
$S$-FasterViT-1 [1, 3, 8, 5] [80, 160, 320, 640] [–, –, 8, 16] 4 41 5.4
$S$-FasterViT-2 [3, 3, 8, 5] [96, 192, 384, 768] [–, –, 12, 24] 4 58 8.8
@@ Appendix E Code
18 self.qkv = $n$.Conv2d(embed_dim, embed_dim*3, 1, bias=True)
19 self.out_proj = $n$.Conv2d(embed_dim, embed_dim, 1, bias=True)